# Algorithms in $4$-manifold topology

### v1

@@ 1. Introduction
  • takes as input two closed, oriented, simply connected, topological $4$-manifolds $X$ and $X^{\prime}$, presented as Kirby diagrams ((((we refer to Section 1.1 for the explanation of what a Kirby diagram in this setting means)))), and
@@ 1.1. Kirby diagrams of simply connected topological 4-manifolds
To encode a closed, oriented, simply connected, topological $4$-manifold $X$ as a finite data set, in Section 3 we introduce the notion of a Kirby diagram of $X$. For this, let $L$ be a framed link in $S^{3}$ with unimodular linking matrix. By attaching $4$-dimensional $2$-handles to $D^{4}$ along $L$, we obtain the diffeomorphism type of a compact, oriented, smooth $4$-manifold $W_{L}$ whose boundary $Y$ is an integral homology $3$-sphere. From Freedman’s work [Fre82] it follows that there exists a contractible, compact, topological $4$-manifold $C$ with boundary $Y$, which we can glue to $W_{L}$ to obtain a topological $4$-manifold $X_{L}$; see Figure 1 for a schematic of the situation. The following theorem shows that $X_{L}$ yields the homeomorphism type of a closed, oriented, simply connected, topological $4$-manifold and that, conversely, any such manifold arises via this construction.
@@ 1.2. The Kirby–Siebenmann invariant
The Kirby-Siebenmann invariant $\operatorname{ks}(X)$ of a topological manifold $X$ (see [FQ90, Section 10.2B] and [KS77, p. 318]) is an element of $H^{4}(X;\mathbb{Z}/2)$. It vanishes if $X$ admits a smooth or a PL structure. Thus, if $X$ is a connected, closed, topological $4$-manifold, the Kirby–Siebenmann invariant is either 00 or $1$. In Section 5 we will use a theorem of Freedman–Kirby (see Theorem 5.3) to show that the Kirby–Siebenmann invariant is computable.
@@ 1.4. Algorithms for the stable classification of 4-manifolds
We say that two smooth $4$-manifolds $X_{1}$ and $X_{2}$ are stably diffeomorphic if there exists a natural number $k$ such that $X_{1}\#_{k}(S^{2}\times S^{2})$ is diffeomorphic to $X_{2}\#_{k}(S^{2}\times S^{2})$. (Note that by [Gom84] two closed, orientable, smooth $4$-manifolds are stably diffeomorphic if and only if they are stably homeomorphic.) The classification of $4$-manifolds up to stable diffeomorphism is coarser, and hence easier to compute, which means that we can leave the realm of simply connected $4$-manifolds. In Section 8 we will use the classification of $4$-manifolds up to stable diffeomorphism to present the following two algorithms.
@@ 1.5. Open questions
A positive resolution of the $11/8$-conjecture (see [GS99, Page 16]) would imply a positive answer to 1.5 for simply connected manifolds.
Question 1.7.
@@
We remark that 1.3 is known to be false in dimensions larger than $4$ [VKF74], whereas the analogues of the other questions seem to be open in all dimensions larger than $4$. Question 1.4 and a theorem similar to Theorem 7.1 for PL and smooth manifolds of dimension $5$ or larger are mentioned in [NW99], building upon the computability of higher homotopy groups [Bro57].
@@ Acknowledgements
We gratefully acknowledge financial support of the SFB 1085 Higher Invariants (Universität Regensburg, funded by the DFG, ID 224262486). Rima Chatterjee was supported by SFB/TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (ID 281071066 - TRR 191). Subhankar Dey was supported by an individual research grant of the DFG (ID 505125645). Daniel Galvin and Paula Truöl would like to thank the Max Planck Institute for Mathematics in Bonn for their hospitality and financial support. Chun-Sheng Hsueh and Naageswaran Manikandan were supported by The Berlin Mathematics Research Center MATH+ funded by the DFG (EXC-2046/1, ID 390685689). Mark Powell was funded by EPSRC New Investigator grant Classifying $4$-manifolds (EP/T028335/2). Martin Tancer was supported by the ERC-CZ project LL2328 of the Ministry of Education of Czech Republic. Simona Veselá was supported by the DFG (EXC-2047/1, ID 390685813).
@@ 2. Intersection forms
where $\operatorname{PD}_{X}\colon H^{2}(X,\partial X;\mathbb{Z})\to H_{2}(X)$ is the Poincaré duality isomorphism given by capping with the fundamental class $[X]\in H_{4}(X,\partial X)$ determined by the orientation. Note that the intersection form is symmetric and bilinear. It follows from the Hurewicz theorem, Poincaré duality, and the universal coefficient theorem that if $X$ is simply connected, then $H_{2}(X)$ is torsion-free. Finally note that it follows from Poincaré duality that the intersection form is unimodular if and only if $X$ is closed or if $\partial X$ is an integral homology $3$-sphere.
@@
    Let $\lambda$ be a symmetric, unimodular, bilinear form over $\mathbb{Z}$ and let $k\in\mathbb{Z}/2$. ((((If $\lambda$ is even, then we assume that $k\equiv\frac{1}{8}\cdot\operatorname{sign}(\lambda)\,\operatorname{mod}2$.)))) There exists a closed, oriented, simply connected, topological $4$-manifold $X$ whose intersection form is isometric to $\lambda$ and with Kirby–Siebenmann invariant $\operatorname{ks}(X)=k\in\mathbb{Z}/2$.
@@ 3. Kirby diagrams of topological 4-manifolds
This framing corresponds to a homotopy class of trivialisations of the normal bundle of $L_{i}$, for each $i$. Each component of $L$ has a closed tubular neighbourhood ${\nu}L_{i}$, and using the framing we obtain the isotopy class of a diffeomorphism from $S^{1}\times D^{2}$ to ${\nu}L_{i}$ (using the standard convention that the 00-framing corresponds to the Seifert framing of $L_{i}$), we refer for example to [GS99] for details.
@@
If the boundary $\partial W_{L}$ is diffeomorphic to #$\#_{m}(S^{1}\times S^{2})$ then we can glue in a copy of $\natural_{m}(S^{1}\times D^{3})$ to obtain a closed, oriented, smooth $4$-manifold $W$. A result by Laudenbach–Poenaru [LP72] implies that the diffeomorphism type of this closed $4$-manifold $W$ is independent of the gluing map and thus $L$ is called Kirby diagram of $W$.
@@ 5. Computing the Kirby–Siebenmann invariant
Freedman–Kirby worked with smooth $4$-manifolds, so for them, the left-hand side was always zero. We will explain below how to deduce the statement we need from statements in the literature. It requires some work because the original references on Rochlin’s theorem were written before Freedman’s work on the disc embedding theorem, so did not consider topological $4$-manifolds. The statement we give was motivated by statements made the introduction to [CST12].
@@
We explain how to obtain the statement from results in the literature. Fix a decomposition $X=W\cup_{Y}C$ as in Section 3, with $W$ smooth, $C$ contractible, and $Y$ an integer homology $3$-sphere. Since, as explained in [FNOP19], the Kirby–Siebenmann invariant is additive under glueing of $4$-manifolds along their boundaries, $\operatorname{ks}(X)=\operatorname{ks}(W)+\operatorname{ks}(C)=\operatorname{ks}(C)$, using that $W$ is smooth in the second equality. It is known that there exists a compact, spin, smooth $4$-manifold $N$ with $\partial N=Y=\partial C$. Then $\operatorname{ks}(N\cup_{Y}C)=\operatorname{ks}(C)$ and $N\cup_{Y}C$ is spin, so
@@
  • outputs a characteristic vector of $V$ with all entries either 00 or $1$.
@@
for some even $(m-k)\times(m-k)$ matrix $E$ over $\mathbb{Z}/2$, i.e. all diagonal entries of $E$ are 00. A characteristic vector of $U$ is given by $c^{T}:=(1,\dots,1,0,\dots,0)$ consisting of $k$-times $1$ followed by $(m-k)$-times 00. To see this, we observe that $x_{i}=x^{2}_{i}$, and the fact that $E$ is even implies that $c^{T}Ux=x^{T}Ux$ for all $x\in(\mathbb{Z}/2)^{m}$. By undoing the basis change we see that $Pc$ is a characteristic vector for $V_{2}$ and thus also for $V$. ∎
@@
We write $X_{L}=W_{L}\cup_{h}C$. The link $L$ induces a preferred isomorphism $H_{2}(X_{L})\cong\mathbb{Z}^{m}$. We use the algorithm from Lemma 4.2 to compute an integer, symmetric, unimodular matrix $V$ representing the intersection form of $X_{L}$ in the above basis. Next, we compute a characteristic vector $c\in H_{2}(X_{L})\cong\mathbb{Z}^{m}$ using the algorithm from Lemma 5.5. As per that lemma, with respect to the standard basis of $\mathbb{Z}^{m}$, the entries of $c$ are either 00 or $1$. Let $L_{c}$ be the associated characteristic sublink of $L$ (i.e. the components of $L$ indexed by the nonzero elements of $c$). Let $K_{c}$ in $S^{3}$ be a knot obtained by some choice of band sum combining the components of $L_{c}$ into a single component.
We claim that $\operatorname{Arf}_{X_{L}}(c)=\operatorname{Arf}(K_{c})$. To see this, we construct a closed, oriented surface $\Sigma$ representing $c$ as follows. Consider the cores of the $2$-handles attached to the components of $L_{c}$. Add the bands used in the construction of $K_{c}$. This gives a disc $D$ in $W_{L}$ with boundary $K_{c}$. Choose an embedded, compact, oriented surface $F$ in the 00-handle $D^{4}$ in $W_{L}$ such that $\partial F=K_{c}$. Take $\Sigma:=F\cup D$ to obtain a closed surface embedded in $W_{L}$ representing $c$. Then $H_{1}(F;\mathbb{Z}/2)\to H_{1}(\Sigma;\mathbb{Z}/2)$ is an isomorphism, and induces an isometry of quadratic forms (because the quadratic enhancements are geometrically computed in exactly the same way [Mat86]). Thus the Arf invariants of these surfaces coincide, and so as claimed
@@ 6. Comparing intersection forms
It remains to discuss the definite case. We will focus here only on the positive definite case, as the proof for the negative definite case can be obtained by replacing $V$ with $-V$.
@@
The strategy to prove Lemma 6.5 is to construct a finite set of lattice points that contains the image of the standard basis of $\mathbb{Z}^{m}$ under any isometry between the two given forms. This reduces the search for an isometry to a finite list of matrices. We start with a lemma from linear algebra for which we introduce the following notation. We write $|x|$ for the Euclidean norm (i.e. $\ell^{2}$-norm) of a vector $x$. For a matrix $A$ with real coefficients we write $\|A\|_{2}$ and $\|A\|_{1}$, respectively, for the operator norms with respect to the $\ell^{2}$ and $\ell^{1}$ norms. We recall that for $A$ an $(m\times m)$ matrix, $\|A\|_{1}=\max_{j}\sum_{i=1}^{m}|a_{ij}|$ and $\|A\|_{2}$ is the square root of the largest eigenvalue of $A^{T}A$. Note that for a symmetric matrix $A$, it follows $\|A\|_{2}=\max\{|\lambda|\,:\,\lambda\,\textrm{ is an eigenvalue of $A$}\}$.
@@
If both matrices are indefinite, we use Corollary 6.4 to determine if $V$ and $V^{\prime}$ are congruent. If both matrices are positive definite we use Lemma 6.5 to determine if $V$ and $V^{\prime}$ are congruent. If both matrices are negative definite, we observe that $-V$ and $-V^{\prime}$ are positive definite and congruent if and only if $V$ and $V^{\prime}$ are congruent, and thus we can reduce this case to the positive definite case. ∎
@@ 8.1. Input
Before stating our main theorems we briefly explain how we can input smooth $4$-manifolds with non-trivial fundamental groups into algorithms. The setup in the previous sections of the paper allows us to input closed, simply connected, topological $4$-manifolds into algorithms. This relies on the theorem saying that every closed, simply connected, topological $4$-manifold can be written as a smooth piece union a contractible topological piece (see Theorem 3.6). This is not known to hold for closed, topological $4$-manifolds with other fundamental groups, so we will have to restrict ourselves to considering smooth $4$-manifolds to be able to easily input manifolds.
@@
  • takes as input an oriented triangulation ((((Kirby diagram)))) of a connected, compact, oriented, smooth $4$-manifold $X$, and
  • outputs a Kirby diagram ((((triangulation)))) of $X$.
@@
It is straightforward to describe algorithms to create out of a Kirby diagram a triangulation. (This is even practically implemented in Regina [Bur24].) On the other hand, since any two triangulations of the same smooth manifold are related by Pachner moves and the set of Kirby diagrams is countable, there exists also an (impractical) algorithm to create a Kirby diagram out of a triangulation. ∎

### v2

@@ 1. Introduction
  • takes as input two closed, oriented, simply connected, topological $4$-manifolds $X$ and $X^{\prime}$, presented as Kirby diagrams ((we refer to Section 1.1 for the explanation of what a Kirby diagram in this setting means)), and
@@ 1.1. Kirby diagrams of simply connected topological 4-manifolds
To encode a closed, oriented, simply connected, topological $4$-manifold $X$ as a finite data set, in Section 3 we introduce the notion of a Kirby diagram of $X$. For this, let $L$ be a framed link in $S^{3}$ with unimodular linking matrix. By attaching $4$-dimensional 22-handles to $D^{4}$ along $L$, we obtain the diffeomorphism type of a compact, oriented, smooth $4$-manifold $W_{L}$ whose boundary $Y$ is an integral homology $3$-sphere. From Freedman’s work [Fre82] it follows that there exists a contractible, compact, topological $4$-manifold $C$ with boundary $Y$, which we can glue to $W_{L}$ to obtain a topological $4$-manifold $X_{L}$; see Figure 1 for a schematic of the situation. The following theorem shows that $X_{L}$ yields the homeomorphism type of a closed, oriented, simply connected, topological $4$-manifold and that, conversely, any such manifold arises via this construction.
@@ 1.2. The Kirby–Siebenmann invariant
The Kirby-Siebenmann invariant $\operatorname{ks}(X)$ of a topological manifold $X$ (see [FQ90, Section 10.2B] and [KS77, p. 318]) is an element of $H^{4}(X;\mathbb{Z}/2)$. It vanishes if $X$ admits a smooth or a PL structure. Thus, if $X$ is a connected, closed, topological $4$-manifold, the Kirby–Siebenmann invariant is either 0 or $1$. In Section 5 we will use a theorem of Freedman–Kirby (see Theorem 5.3) to show that the Kirby–Siebenmann invariant is computable.
@@ 1.4. Algorithms for the stable classification of 4-manifolds
We say that two smooth $4$-manifolds $X_{1}$ and $X_{2}$ are stably diffeomorphic if there exist natural numbers $k_{1}$ and $k_{2}$ such that $X_{1}\#_{k_{1}}(S^{2}\times S^{2})$ is diffeomorphic to $X_{2}\#_{k_{2}}(S^{2}\times S^{2})$. Note that by [Gom84] two closed, orientable, smooth $4$-manifolds are stably diffeomorphic if and only if they are stably homeomorphic. The classification of $4$-manifolds up to stable diffeomorphism is coarser, and hence easier to compute, which means that we can leave the realm of simply connected $4$-manifolds. In Section 8 we will use the classification of $4$-manifolds up to stable diffeomorphism to present the following two algorithms.
@@
Remark 1.3.
In the literature one also finds stable diffeomorphism defined with $k_{1}=k_{2}$. We could have also used this definition, but then would have had to require that $\chi(X_{1})=\chi(X_{2})$. Since the Euler characteristic is straightforward to algorithmically compute from a triangulation, Theorems 8.3 and 8.4 also hold with the more rigid alternative definition of stable diffeomorphism.
@@ 1.5. Open questions
A positive resolution of the $11/8$-conjecture (see [GS99, Page 16]) would imply a positive answer to Question 1.6 for simply connected manifolds.
Question 1.8.
@@
We remark that Question 1.4 is known to be false in dimensions larger than $4$ [VKF74], whereas the analogues of Questions 1.6 and 1.8 seem to be open in all dimensions larger than $4$. Question 1.5 and a theorem similar to Theorem 7.1 for PL and smooth manifolds of dimension $5$ or larger are mentioned in [NW99], building upon the computability of higher homotopy groups [Bro57].
@@ Acknowledgements
We gratefully acknowledge financial support of the SFB 1085 Higher Invariants (Universität Regensburg, funded by the DFG, ID 224262486). Rima Chatterjee was supported by SFB/TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (ID 281071066 - TRR 191). Subhankar Dey was supported by an individual research grant of the DFG (ID 505125645). Daniel Galvin and Paula Truöl would like to thank the Max Planck Institute for Mathematics in Bonn for their hospitality and financial support. Alejandro García Rivas was supported by MSCA-DN CaLiForNIA - 101119552. Chun-Sheng Hsueh and Naageswaran Manikandan were supported by The Berlin Mathematics Research Center MATH+ funded by the DFG (EXC-2046/1, ID 390685689). Marc Kegel is supported by the DFG, German Research Foundation, (Project: 561898308); by a Ramón y Cajal grant (RYC2023-043251-I) and the project PID2024-157173NB-I00 funded by MCIN/AEI/10.13039/501100011033, ESF+ and FEDER, EU; and by a VII Plan Propio de Investigación y Transferencia (SOL2025-36103) of the University of Sevilla. Mark Powell was funded by EPSRC New Investigator grant Classifying $4$-manifolds (EP/T028335/2). Martin Tancer was supported by the ERC-CZ project LL2328 of the Ministry of Education of Czech Republic. Simona Veselá was supported by the DFG (EXC-2047/1, ID 390685813).
Open access
For the purpose of open access, the authors have applied a CCBY license to this version, and will do the same to any author-accepted manuscript arising from this submission.
@@ 2. Intersection forms
where $\operatorname{PD}_{X}\colon H^{2}(X,\partial X;\mathbb{Z})\to H_{2}(X)$ is the Poincaré duality isomorphism given by capping with the fundamental class $X]\in H_{4}(X,\partial X)$ determined by the orientation. Note that the intersection form is symmetric and bilinear. It follows from the Hurewicz theorem, Poincaré duality, and the universal coefficient theorem that if $X$ is simply connected, then $H_{2}(X)$ is torsion-free. Finally note that it follows from Poincaré duality that the intersection form is unimodular if and only if $X$ is closed or if $\partial X$ is an integral homology $3$-sphere.
@@
    Let $\lambda$ be a symmetric, unimodular, bilinear form over $\mathbb{Z}$ and let $k\in\mathbb{Z}/2$. ((If $\lambda$ is even, then we assume that $k\equiv\frac{1}{8}\cdot\operatorname{sign}(\lambda)\,\operatorname{mod}2$.)) There exists a closed, oriented, simply connected, topological $4$-manifold $X$ whose intersection form is isometric to $\lambda$ and with Kirby–Siebenmann invariant $\operatorname{ks}(X)=k\in\mathbb{Z}/2$.
@@ 3. Kirby diagrams of topological 4-manifolds
This framing corresponds to a homotopy class of trivialisations of the normal bundle of $L_{i}$, for each $i$. Each component of $L$ has a closed tubular neighbourhood ${\nu}L_{i}$, and using the framing we obtain the isotopy class of a diffeomorphism from $S^{1}\times D^{2}$ to ${\nu}L_{i}$ (using the standard convention that the 0-framing corresponds to the Seifert framing of $L_{i}$), we refer for example to [GS99] for details.
@@
If the boundary $\partial W_{L}$ is diffeomorphic to $_{m}(S^{1}\times S^{2})$ then we can glue in a copy of $\natural_{m}(S^{1}\times D^{3})$ to obtain a closed, oriented, smooth $4$-manifold $W$. A result by Laudenbach–Poenaru [LP72] implies that the diffeomorphism type of this closed $4$-manifold $W$ is independent of the gluing map and thus $L$ is called Kirby diagram of $W$.
@@ 5. Computing the Kirby–Siebenmann invariant
Freedman–Kirby worked with smooth $4$-manifolds, so for them, the left-hand side was always zero. We will explain below how to deduce the statement we need from statements in the literature. It requires some work because the original references on Rochlin’s theorem were written before Freedman’s work on the disc embedding theorem, so did not consider topological $4$-manifolds. The statement we give was motivated by statements made in the introduction to [CST12].
@@
We explain how to obtain the statement from results in the literature. Fix a decomposition $X=W\cup_{Y}C$ as in Section 3, with $W$ smooth, $C$ contractible, and $Y$ an integer homology $3$-sphere. Since, as explained in [FNOP19], the Kirby–Siebenmann invariant is additive under glueing of $4$-manifolds along their boundaries, $\operatorname{ks}(X)=\operatorname{ks}(W)+\operatorname{ks}(C)=\operatorname{ks}(C)$, using that $W$ is smooth in the second equality. It is known that there exists a compact, spin, smooth $4$-manifold NN with $\partial N=Y=\partial C$. Then $\operatorname{ks}(N\cup_{Y}C)=\operatorname{ks}(C)$ and $N\cup_{Y}C$ is spin, so
@@
  • outputs a characteristic vector of $V$ with all entries either 0 or $1$.
@@
for some even $(m-k)\times(m-k)$ matrix $E$ over $\mathbb{Z}/2$, i.e. all diagonal entries of $E$ are 0. A characteristic vector of $U$ is given by $c^{T}:=(1,\dots,1,0,\dots,0)$ consisting of $k$-times 11 followed by (mk)(m-k)-times 0. To see this, we observe that $x_{i}=x^{2}_{i}$, and the fact that $E$ is even implies that $c^{T}Ux=x^{T}Ux$ for all $x\in(\mathbb{Z}/2)^{m}$. By undoing the basis change we see that $Pc$ is a characteristic vector for $V_{2}$ and thus also for $V$. ∎
@@
We write $X_{L}=W_{L}\cup_{h}C$. The link $L$ induces a preferred isomorphism $H_{2}(X_{L})\cong\mathbb{Z}^{m}$. We use the algorithm from Lemma 4.2 to compute an integer, symmetric, unimodular matrix $V$ representing the intersection form of $X_{L}$ in the above basis. Next, we compute a characteristic vector $c\in H_{2}(X_{L})\cong\mathbb{Z}^{m}$ using the algorithm from Lemma 5.5. As per that lemma, with respect to the standard basis of $\mathbb{Z}^{m}$, the entries of $c$ are either 0 or $1$. Let $L_{c}$ be the associated characteristic sublink of $L$ (i.e. the components of $L$ indexed by the nonzero elements of $c$). Let $K_{c}$ in $S^{3}$ be a knot obtained by some choice of band sum combining the components of $L_{c}$ into a single component.
We claim that $\operatorname{Arf}_{X_{L}}(c)=\operatorname{Arf}(K_{c})$. To see this, we construct a closed, oriented surface $\Sigma$ representing $c$ as follows. Consider the cores of the $2$-handles attached to the components of $L_{c}$. Add the bands used in the construction of $K_{c}$. This gives a disc $D$ in $W_{L}$ with boundary $K_{c}$. Choose an embedded, compact, oriented surface $F$ in the 0-handle $D^{4}$ in $W_{L}$ such that $\partial F=K_{c}$. Take $\Sigma:=F\cup D$ to obtain a closed surface embedded in $W_{L}$ representing $c$. Then $H_{1}(F;\mathbb{Z}/2)\to H_{1}(\Sigma;\mathbb{Z}/2)$ is an isomorphism, and induces an isometry of quadratic forms (because the quadratic enhancements are geometrically computed in exactly the same way [Mat86]). Thus the Arf invariants of these surfaces coincide, and so as claimed
@@ 6. Comparing intersection forms
It remains to discuss the definite case. We will focus here only on the positive definite case, as the proof for the negative definite case can be obtained by replacing $V$ with V-V.
@@
[…]
@@
If both matrices are indefinite, we use Corollary 6.4 to determine if $V$ and $V^{\prime}$ are congruent. If both matrices are positive definite we use Lemma 6.5 to determine if $V$ and $V^{\prime}$ are congruent. If both matrices are negative definite, we observe that V-V and $V^{\prime}$ are positive definite and congruent if and only if $V$ and $V^{\prime}$ are congruent, and thus we can reduce this case to the positive definite case. ∎
@@ 8.1. Input
Before stating our main theorems we briefly explain how we can input smooth $4$-manifolds with non-trivial fundamental groups into algorithms. The setup in the previous sections of the paper allows us to input closed, simply connected, topological $4$-manifolds into algorithms. This relies on the theorem saying that every closed, simply connected, topological $4$-manifold can be written as a smooth piece union a contractible topological piece (see Theorem 3.6). This is not known to hold for closed, topological $4$-manifolds with other fundamental groups, but is known to hold stably, i.e. every topological 4-manifold is stably homeomorphic to a smooth manifold union a contractible piece. [^2] If the manifold $X$ is stably smoothable then the statement is clear. If it is not consider $X\#\overline{E_{8}}\#E_{8}$, which is stably homeomorphic to $X$. Now $E_{8}$ is smooth away from a contractible piece $C$, and so we can stably smooth $X\#\overline{E_{8}}\#E_{8}$ away from $C$. Likely one could combine the work of the previous sections and this one to obtain results about stable homeomorphism for general topological 4-manifolds, but this would require us to input smooth data that is only abstractly known to emerge from stabilisations. Since this is somewhat unsatisfactory, we will instead only consider smooth 4-manifolds as inputs.
@@
  • takes as input an oriented triangulation ((Kirby diagram)) of a closed, oriented, connected, smooth $4$-manifold $X$, and
  • outputs a Kirby diagram ((triangulation)) of $X$.
@@
It is straightforward to describe algorithms to create a triangulation out of a Kirby diagram. (This is even practically implemented in Regina [Bur24].) On the other hand, since any two triangulations of the same smooth manifold are related by Pachner moves and the set of Kirby diagrams is countable, there also exists an (impractical) algorithm to create a Kirby diagram out of a triangulation. ∎